# Sampling reduced density matrix to extract fine levels of entanglement spectrum

### v2

@@ Abstract
Conclusions and outlooks.- We propose a practical scheme to extract the fine levels of entanglement spectrum from the method of quantum Monte Carlo simulation combined with exact diagonalization. It can extract the fine information of low-lying ES for large-scale quantum many-body systems whose environment has huge degrees of freedom. Using this method, we answer the previous contradiction for the entanglement spectrum in a Heisenberg ladder system. Furthermore, we calculate the ES of a 2D AFM Heisenberg model within two different cut-geometries. Both the cornered and cornerless cases show the TOS character which reflects the continuous symmetry breaking here.
@@
In summary, the reduced density matrix can be obtained in larger size via QMC within this frame. Therefore, other observables such as von Neumann entropy and off-diagonal measurements can also be extracted in this way.
@@
Acknowledgement.- We would like to thank Wei Zhu, Yin Tang, Xue-Feng Zhang, Shangqiang Ning and Han-Qing Wu for fruitful discussions. We acknowledge the start-up funding of Westlake University. The authors also acknowledge Beijng PARATERA Tech Co.,Ltd.(https://www.paratera.com/) for providing HPC resources that have contributed to the research results reported within this paper.

### v3

@@ Abstract
Conclusions and outlooks.- We propose a practical scheme to extract the fine levels of entanglement spectrum from the method of quantum Monte Carlo simulation combined with exact diagonalization. It can extract the fine information of low-lying ES for large-scale quantum many-body systems whose environment has huge degrees of freedom. Using this method, we explain the previous discrepancy for the entanglement spectrum in a Heisenberg ladder system. Furthermore, we calculate the ES of a 2D AFM Heisenberg model within two different cut-geometries. Both the cornered and cornerless cases show the TOS character which reflects the continuous symmetry breaking here.
@@
In summary, the reduced density matrix can be obtained in larger size via QMC within this simple frame. Therefore, other observables such as von Neumann entropy and off-diagonal measurements can also be extracted in this way. This low-technical-barrier algorithm can be widely used in quantum Monte Carlo simulations and is easy to be parallelized on computer.
@@
Acknowledgement.- We would like to thank Wei Zhu, Yin Tang, Chen Cheng, Xue-Feng Zhang, Shangqiang Ning, Marcello Dalmonte and Han-Qing Wu for fruitful discussions. We acknowledge the start-up funding of Westlake University. The authors also acknowledge the HPC centre of Westlake University and Beijng PARATERA Tech Co.,Ltd. for providing HPC resources that have contributed to the research results reported within this paper.